OpenStudy (anonymous):

integrate by parts: ∫xln(1+x)dx

OpenStudy (anonymous):

Integrating by part will do, i think.

OpenStudy (anonymous):

\[\large \int u \, dv=uv-\int v \, du.\!\] Set \(u\) to something easily differentiable.

OpenStudy (anonymous):

u=x dv=ln(1+x)

OpenStudy (anonymous):

i got